OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? Use the following System of Equations to answer the question: y = x – 3 y = 5 – 3x What is the value of x? a. 1 b. –1 c. 2 d. –2

OpenStudy (anonymous):

giv me one sec

OpenStudy (igreen):

We write: x - 3 = 5 - 3x Add 3 to both sides: x = 8 - 3x Add 3x to both sides: 4x = 8 Now can you divide 8 / 4? @ninaesb

OpenStudy (anonymous):

2 @iGreen

OpenStudy (igreen):

Yes, that's your answer.

OpenStudy (damonte212):

yah i agree its C.